OpenStudy (anonymous):

I cant remember the formula for this can anyone help? Perimeter of a rectangle is 600 yards. What are the dimensions of the rectangle if the length is 80 yards more than the width? thanks :)

OpenStudy (anonymous):

Let the width be x. So length = x + 80 2(l + b) = 600 2(x + 80 + x) = 600 Can you solve now?

OpenStudy (anonymous):

p=2*(l+b)=600 l=b+80 solve these two equations

OpenStudy (anonymous):

I can solve it but what does b stand for?

OpenStudy (anonymous):

b is breadth or width.